OpenStudy (anonymous):

How do u solve (2/p)^-3

OpenStudy (superdavesuper):

Does it ask for a simplified form?

OpenStudy (anonymous):

Yea

OpenStudy (superdavesuper):

okay then (2/p)^-3=(2^-3)/(p^-3) then u can simplify it further...

OpenStudy (jdoe0001):

\(\bf a^{-{\color{red} n}} = \cfrac{1}{a^{\color{red} n}}\\ \quad \\ \quad \\ \left(\cfrac{2}{p}\right)^{-3}\implies \cfrac{2^{-3}}{p^{-3}}\implies \cfrac{\frac{1}{\square ?}}{\frac{1}{\square ?}}\\ \quad \\ \cfrac{1}{\square ?}\cdot \cfrac{\square ?}{1}\)

OpenStudy (anonymous):

Thank you